# SoMeR: Multi-View User Representation Learning for Social Media

### v1

@@ Abstract
User representation learning aims to capture user preferences, interests, and behaviors in low-dimensional vector representations. These representations have widespread applications in recommendation systems and advertising; however, existing methods typically rely on specific features like text content, activity patterns, or platform metadata, failing to holistically model user behavior across different modalities. To address this limitation, we propose SoMeR, a Social Media user Representation learning framework that incorporates temporal activities, text content, profile information, and network interactions to learn comprehensive user portraits. SoMeR encodes user post streams as sequences of timestamped textual features, uses transformers to embed this along with profile data, and jointly trains with link prediction and contrastive learning objectives to capture user similarity. We demonstrate SoMeR’s versatility through two applications: 1) Identifying inauthentic accounts involved in coordinated influence operations by detecting users posting similar content simultaneously, and 2) Measuring increased polarization in online discussions after major events by quantifying how users with different beliefs moved farther apart in the embedding space. SoMeR’s ability to holistically model users enables new solutions to important problems around disinformation, societal tensions, and online behavior understanding.
[…]
123
1 Introduction
User representation learning aims to learn low-dimensional vector representations of users that capture their preferences, interests, and behaviors [22]. These representations are used widely in commercial applications, such as personalized recommendation, targeted advertising, and user modeling [2, 42, 5, 17, 37]. Applications of user representation learning extend well beyond commercial sector to the social media domain, where they are used to uncover latent factors of user online behavior that give insights into public attitudes and societal trends, and help understand the formation of online echo chambers [29]. Prior studies have utilized learned representations to detect social bots and inauthentic information operation drivers [26], identify suicidal ideation [36] and detect hate speech [31, 8]. These methods, however, usually depend on specific features, for example, content-based features that use text [14] or combine text with images [29], activity-based features [25, 26], or platform-specific features [1, 7]. As a result, these approaches are not able to holistically model user behavior by combining multiple types features from content, temporal activity and interactions between users.
@@ 1 Introduction
Multi-view user representation learning for social media that combines multiple streams of evidence poses a number of challenges. User populations on social media are highly heterogeneous, composed of individuals with different beliefs, attitudes, interactions and behaviors. A few users within a population are prolific posters while the vast majority of others post only infrequently. As a result, the temporal user activity is sparse, which presents a challenge for traditional time series analysis methods. Similarly, a few users have large numbers of followers while the vast majority are followed by a handful of others. In addition, for many tasks, ground truth data is not available, or it is difficult to obtain, which makes it hard to train machine classifiers.
@@
To address these challenges, we propose SoMeR, a Social Media user Representation Learning framework, which incorporates 1) temporal activities, 2) texts of posts, 3) profile information and 4) network interactions to learn a comprehensive portrait of online users. These features are universal across different social platforms, making our framework very flexible and adaptive. Moreover, this framework allows us to discover similar users in heterogeneous populations with different beliefs, attitudes, and behaviors, thereby creating new solutions to difficult and important problems.
@@
The method works as follows. We first encode the stream of a user’s posts as a sequence of triplets composed of $(timestamp,textualfeature,value)$. This helps address the challenge of modeling sparse temporal activity data where users post only infrequently. We encode the contextual information of these triplets into an embedding using a transformer-based architecture [40]. We combine this triplet embedding with user profile embedding, and impose two jointly trained objectives: (1) network link prediction to learn interactions between users, and (2) contrastive learning to pull users with similar posting histories closer and push dissimilar users farther away. In the end, the model learns an embedding space that is aware of user similarity and heterogeneity across temporal, textual, network connection and user profile dimensions.
The pre-training step described above learns enriched user representations from multiple features, which can be used in unsupervised settings where annotated data is hard to obtain. The method can also be used within supervised learning. By fine-tuning our model with fewer data, we can adapt them to a range of downstream tasks. For example, by fine-tuning with user ideology, we create an ideologically-aware embedding space to answer questions about polarization.
The framework is extremely versatile and can be adapted to study diverse research questions. We demonstrate framework’s generalizability and versatility in two applications. The first application allows us to identify inauthentic accounts associated with online information operations from among a multitude of online activity, i.e., information operation (IO) drivers. Social media has emerged as a battleground for IOs, enabling malicious actors to mount coordinated influence campaigns to amplify societal divisions or sway public opinion [44]. The proposed method achieves this goal by identifying accounts that post similar texts at similar times — a signature of coordination. We learn user representations from social media data and train a supervised model to detect IO drivers that share similar content at similar times, which is a signature of coordinated online influence campaigns [28, 26]. When test on known IOs from three countries that have been verified by the platform X (formerly Twitter), our method shows outstanding performance in detecting covert influence campaigns.
Our second application uses learned representations to measure how events widen partisan division and increase polarization of online discussions. Specifically, using data from online discussions about the overturning of federal protections for abortion in the U.S., we show the impact of the decision on the activities of partisans. We use the proposed user representation method to study how populations with different political beliefs responded to the decision. The analysis reveals an increased polarization: users with the same political belief moved closer, whereas users with different believes moved farther apart. These two applications highlight the flexibility and universality of the proposed method.
Our proposed framework helps to bridge the divide between user representation learning and socio-political analysis, providing a promising method to deepen our understanding of user heterogeneity and its dynamics, and paving the way for more informed decisions and interventions in an increasingly interconnected world.
2 Related Works
User Representation Learning
User representation learning has gained widespread interest in the recommendation system research area due to its ability to capture meaningful and compact embeddings that represent important characteristics of users’ behaviors and preferences [45]. Over the years, researchers have developed many methods [22], including matrix and tensor factorization [2, 15] and deep learning based models, such as auto-encoders [42, 47] and recently more sophisticated transformer-based architectures [5, 17, 37]. Researchers have also advanced the training techniques for these deep learning methods, such as incorporating contrastive learning to learn similarity in data without the need of human labels [27, 5]. Improving from task-specific methods [18, 12] in the early days, many works have also contributed to build universal user representation learning methods which can be generalized to different downstream tasks [45, 17, 37]. The demand for recommendation systems has largely advanced the techniques and methods in user representation learning. However, many of these works formulate the major task to be customization and adaptation of systems to the user’s specific needs [22], and most of the methods have only been tested on product preference prediction or user profiling.
User Understanding in Social Domains
Beyond recommendation systems, user representation learning is also useful to understand online users and communities through their behaviors and opinions [29]. Mueen et al. [25] and Nwala et al. [26] uses temporal activity features to detect online bot and IO drivers (sometimes known as coordinated users); Hallac et al. [14, 13] experimented with different textual embedding methods, such as TF-IDF, doc2vec and BERT for social media user representation learning; Perozzi et al. [30] uses network features and node prediction to learn user interests; Wang et al. [43] use network embedding method to learn about communities; and Ribeiro et al. [34] use both text and network features to detect hate speech and hateful users online. However, there lacks a universal framework for social analysis that is generalizable to different downstream tasks and incorporates all of the textual, temporal, profile and network features. Our framework, SoMeR, addresses these gaps.
3 Methods
@@
We propose a self-supervised framework to learn a latent user embedding space based on the architecture shown in Figure 1. From each user’s history, a timeline of texts, we first extract certain textual features, such as the sentence embeddings. Next, to better learn from users with sparse activities, we format the textual features and timestamps into triplets of observations $(timestamp,feature,value)$. These triplets pass through a Triplet Encoder, a transformer-based contextual learning module, and a fusion attention layer, being encoded into a user history embedding. We concatenate it with user’s profile embedding from a separate module, obtaining a complete user embedding. We train these encoding modules with two self-supervised objectives: network link prediction that learns patterns of interactions, including sharing, following or other connections, and contrastive loss that learns user similarity with respect to posting history. This method learns user similarity in a heterogeneous user population without the need for time-consuming human annotations. The method can be easily adapted for various downstream tasks such as supervised learning and unsupervised similarity search.
@@
In this study, we consider each user’s history to be a collection of timestamped texts. Each text can be an original post, a repost, a reply, etc. There can be other types of user activity, such as likes or views, that also provide valuable information. We plan to incorporate these data in future works. To learn from a user’s history, first we need to extract the desired textual features that are appropriate for the downstream task. These features can be topics discussed in the posts, emotions expressed in them, or contextual features extracted as text embeddings [11]. In our experiments, we find that using contextual sentence-BERT embeddings [33] leads to a flexible and powerful representation, and hence use this approach. However, BERT embeddings yield more complex features than other methods, which can slow performance. To reduce model size and complexity, we perform dimensionality reduction using principal component analysis (PCA) to reduce BERT embeddings to the first five components and treat them as textual features.
@@
Social media users exhibit highly diverse posting behaviors. A small fraction of users generate most of the contents, whereas majority of users only have few posts over a long period of time [21]. This leads to a sparse matrix of multivariate time series as the input data, which makes representation learning harder. We take inspiration from [38], and represent user posting history as a collection of triplets $(timestamp,feature,value)$. Thus, a dataset of $M$ users can be represented as $U$:
@@
Each user $u$ is characterized by their profile feature vector $d^{u}\in\mathbb{R}^{D}$ and their posting history $\mathbf{X}^{u}$, which is a set of $N$ triplets including the timestamp $t_{n}$, the feature category $f_{n}$ and the value of this feature category $v_{n}$. Note that there can be more than one feature-value pairs at one time point, and therefore $N$ does not necessarily equal to the total time points.
Tipirneni and Reddy [38] shows the effectiveness of using a feed-forward network to embed continuous values. We therefore use two separate feed-forward networks to encode timestamp $t_{n}$ and value $v_{n}$ into time embedding $e^{t}_{n}\in\mathbb{R}^{K}$ and value embedding $e^{v}_{n}\in\mathbb{R}^{K}$ respectively, where $K$ is the hidden dimension of these embeddings. These networks have one linear layer followed by a tanh activation function. For the feature categories, we use a lookup-table encoder to generate a feature embedding $e^{f}_{n}\in\mathbb{R}^{K}$. Lastly, we add up these three embeddings to be the triplet embedding $e_{n}^{triplet}\in\mathbb{R}^{K}$.
@@
The transformer architecture has been shown to have great performance in representation learning for time series and user behavior sequences [46, 37]. Therefore we choose to use the transformer encoder to better extract a latent representation from the triplet embedding $e_{n}^{triplet}\in\mathbb{R}^{K}$. We use $L$ transformer layers. Each layer has $H$ attention heads with learnable key, query and value weights $W^{k},W^{q},W^{v}\in\mathbb{R}^{K\times P}$ where $P$ is the hidden dimension size for these weights. After all attention heads are added up and layer normalized, the embedding vector is projected back to $K$-dimension through a feed-forward network. This network includes two layers with hidden dimension $2K$ and a ReLU activation in the middle. Lastly, layer normalization is applied again. From this transformer module we obtain $e_{n}^{trans}\in\mathbb{R}^{K}$. In our experiments and applications, data sizes are within 20 million, and therefore we choose to have a small transformer module with $L=2$, $H=4$ and $P=K/H$ which we show in § 4 and § 5 works well for our purposes.
@@
Individuals connected in social networks tend to be similar, and their behaviors are often affected by other users in the network [24]. Therefore it’s crucial to include network connection features when learning user representations [1]. We design a self-supervised network link prediction objective to train our model to learn interaction activities such as sharing, following and commenting. It is a feed-forward module to perform link prediction, with a binary cross-entropy loss. During training, we consider all pairs of distinct users in each batch. The feature for link prediction is the concatenated embedding of a user pair, and we obtain the links from a self-defined network as the binary labels (e.g. whether a user repost another user). This is described as following:
@@
Contrastive Learning and Data Augmentation
Contrastive learning aims to obtain a latent embedding space in which similar samples are closer and distinct samples are farther from each other. Many prior works in user representation learning have shown its success [5, 37]. We adopt self-supervised contrastive learning to learn user similarity in their posting histories. The InfoNCE [27] loss function uses categorical cross-entropy loss to optimize the negative log probability of classifying one positive or similar sample correctly among a set of negative or unrelated samples. In our case, it can be written as:
@@
To generate the positive sample paired to each user history embedding, we perform data augmentation on users’ triplet data. Although researchers have proposed many time series data augmentation methods [44], many classical methods are not applicable to our scenario. For example, a user history cannot be sub-sequenced or shuffled in time domain. We use another efficient and simple way - bootstrapping with replacement from existing triplets $\{(t_{n},f_{n},v_{n})\}_{n=1}^{N}$. This ensures the generated set of triplets is similar. We incorporate noise into augmentation by
@@
During training, we tune $\gamma$ by grid searching in $[0.1,0.5,1,2,5]$. The results of the two applications below shows the effectiveness of this augmentation method.
@@
where $\lambda$ is a hyperparamter to balance between two losses. We perform grid search on $\lambda$ in $[0.1,1,5,10]$ and decide to use $\lambda=1$. We also tune the hidden dimension $K$ used in different modules with a grid search in $[32,64,128]$ and decide to use $K=64$. We use Adam optimizer, a learning rate of 5e-5, a cosine decay learning scheduler, an early stopping mechanism monitored by overall loss function, a maximum epoch of 60 and a batch size of 128.
@@
To verify that our self-supervised framework indeed learns temporal activity and textual features from the triplet data, we test it on two synthetic datasets. We use synthesized numerical features to mimic textual features in real data. One dataset consists of clusters that vary in timestamps but have the same features and values, and another dataset consists of clusters that vary only in features and values, but having all the same timestamps. On both datasets, our model can successfully detect the clusters, indicating it is able to learn heterogeneity in both temporal activities and numerical feature values (see S.I. for details).
In the next sections, we use two applications - IO driver detection and political polarization analysis - to illustrate the effectiveness and versatility of our framework.
4 Detecting Drivers of Information Operations
Information operations use strategically organized efforts to sway and manipulate public opinions, contaminating the online information ecosystem with disinformation. The rapidly expanding social media platforms provide fertile grounds for these operations [28]. Therefore, it is a critical to detect these IO drivers. IO drivers have different tactics from each other and from non-IO accounts thus leading to distinct behaviors we can identify, such as (1) co-sharing the same posts or URLs, (2) using an identical sequence of hashtags, (3) synchronized behaviors, and (4) high textual similarity in posts. Researchers have developed some effective methods for IO drivers using these features [28, 26, 23].
However, most of these methods are specifically designed for the X platform. As many other platforms, such as Reddit, also host suspicious operations [35], there is a need to develop more generalizable methods for coordination detection. Moreover, IO driver behavior may not be identical to another account, but their behavior could nonetheless be distinguishable. Our user representation learning framework can account for temporal, textual, profile and network connection features, within which synchronized behaviors are ingrained. Moreover, our framework is not limited to any platform-specific features. Here we perform supervised fine-tuning on top of our framework to build an IO driver detector, and test it on three different campaigns to demonstrate its efficacy.
@@ 4 Detecting Drivers of Information Operations
We evaluate our method on an X dataset that prior IO driver detection methods have benchmarked on [26, 23]. This set of IO drivers from multiple operations in 21 countries were suspended and released by X because they were associated with malicious IOs and violated the platform terms. We select four campaigns — one in China involving a large number of accounts, one small operation in Egypt and UAE, and two operations in Venezuela which are combined into one dataset for testing the multi-campaign scenario. In future works, we plan to expand our framework to other social platforms.
We collect a set of control users by first collecting the top five keywords and top five hashtags for each IO driver within our X datasets. We then extract 10 random posts that were posted within the timeframe that the IO driver was active (between their first and last post). For each post, we find the post author and query all their posts made within the timeframe the IO driver is active. Table 1 shows the information about these campaigns. We split data by 70%-20%-10% for training, validation and testing.

### v2

# SoMeR: A Multi-View Social Media User Representation Learning Framework

@@ Abstract
Social media user representation learning aims to capture user preferences, interests, and behaviors in low-dimensional vector representations. These representations are critical to a range of social problems, including predicting user behaviors and detecting inauthentic accounts. However, existing methods are either designed for commercial applications, or rely on specific features like text contents, activity patterns, or platform metadata, failing to holistically model user behavior across different modalities. To address these limitations, we propose SoMeR, a Social Media user Representation learning framework that incorporates temporal activities, text contents, profile information, and network interactions to learn comprehensive user portraits. SoMeR encodes user post streams as sequences of time-stamped textual features, uses transformers to embed this along with profile data, and jointly trains with link prediction and contrastive learning objectives to capture user similarity. We demonstrate SoMeR’s versatility through three applications: 1) Identifying information operation driver accounts, 2) Measuring online polarization after major events, and 3) Predicting future user participation in Reddit hate communities. SoMeR provides new solutions to better understand user behavior in the socio-political domains, enabling more informed decisions and interventions.
Introduction
Understanding and modeling user behavior on social media is a critical challenge for platforms, researchers, and society at large. As social media continues to shape public discourse, civic processes, and responses to disasters, the ability to accurately represent user preferences, opinions, and behavioral patterns has become increasingly urgent. These user representations serve as foundational building blocks for numerous downstream applications, from personalized content recommendation and detecting social bots (Nwala et al. 2023) to identifying suicidal ideation (Sawhney et al. 2021) and hate speech (Qian et al. 2018; Del Tredici et al. 2019). The complexity of this challenge stems from the multi-faceted nature of user behavior—individuals express themselves through various modalities including text posts, sharing patterns, temporal rhythms of activity, and social connections, all of which evolve dynamically over time.
To address this problem, researchers have developed methods for user representation learning (URL), which aim to learn low-dimensional vector representations that capture user preferences, interests, and behaviors (Li and Zhao 2021). These representations are popular in commercial applications, such as personalized recommendations and targeted advertising (Bhargava et al. 2015; Wang et al. 2015; Hou et al. 2022; Shin et al. 2023). Applications of URL extend well beyond the commercial sector to the social and political domains, where they are used to uncover latent factors of online behavior that give insights into public attitudes and societal trends (Pan and Ding 2019).
However, most advanced URL approaches, designed for recommendation and built on commercially relevant features (Li and Zhao 2021; Purificato et al. 2024), are not well-suited for social media. In addition, the current URL methods developed for social media usually depend on a subset of user features, such as text (Hallac et al. 2021), images (Pan and Ding 2019), activity (Nwala et al. 2023), or platform-specific features (AlMahmoud and AlKhalifa 2018; Dahiya et al. 2021). Multi-view approaches exist but are restricted to combining only text and network features, and many remain task- or platform-specific (Feng et al. 2021; Shen et al. 2023). As a result, there is a scarcity of generalized approaches for socio-political problems that (1) holistically model user behavior by combining multiple features from content, temporal activity and user interactions, and (2) can be flexibly applied to different research scenarios.
@@ Introduction
Creating universal multi-view user representation learning that combines multiple features poses a number of challenges. The social media user populations are highly heterogeneous, composed of individuals with different beliefs, attitudes, interactions, and behaviors. A few users within a population are prolific posters while the vast majority are not. As a result, the temporal user activity is sparse, which presents a challenge when incorporating activity features into a representation. Moreover, the number of connections varies dramatically between users and also changes from time to time, which poses challenges to incorporating network information. Finally, ground truth data to train classifiers is often difficult to obtain, which makes a supervised model representation challenging.
@@
To address these challenges, we propose SoMeR, Social Media user Representation learning framework, which incorporates 1) temporal activities, 2) post text, 3) profile information and 4) network interactions to learn a comprehensive portrait of online users. These features are universal across different social platforms, making our framework flexible and adaptable. This framework allows us to discover similar users in populations with heterogeneous beliefs, attitudes, and behaviors.
@@
The method works as follows. We first encode user posts as a sequence of triplets of the form (timestamp, textual feature, value), which augments typically limited time series data by incorporating a variety of features from each post. We encode the contextual information of these triplets into an embedding using a transformer-based architecture (Vaswani et al. 2017). We combine this triplet embedding with a user profile embedding, and impose two jointly trained objectives: (1) network link prediction to learn interactions between users, and (2) contrastive learning to pull similar users closer and push dissimilar users farther away. As a result, our model learns an embedding space that is aware of user similarity and heterogeneity across temporal, textual, network connection, and user profile dimensions. This pre-training step can be used in unsupervised settings where annotated data is hard to obtain. Moreover, by fine-tuning our model on a few supervision data samples, we can adapt it to various downstream tasks.
We demonstrate the generalizability and versatility of the framework by studying three novel and diverse research questions on two different social media platforms. We first identify Information Operation (IO) drivers, which are malicious accounts used to amplify societal divisions or sway public opinion (Wen et al. 2020). We train a SoMeR-based model on a novel ground truth information operations dataset and detect IO drivers with accuracy that matches or exceeds baselines. We next use SoMeR to measure how events deepen polarization. We analyze discussions on X about the US Supreme Court (SCOTUS) ruling overturning federal abortion protections, and find that the ruling increased partisan differences in text and activity of users. Finally, we predict future user participation in Reddit hate communities (subreddits). Prior research shows a direct link between online hate and offline attacks (Müller and Schwarz 2021), thus detecting who will become active in hate communities can improve interventions to reduce offline hate. SoMeR achieves unexpectedly high accuracy, even if users do not post months before posting their first hate group.
Our framework advances the state-of-the-art by filling gaps in generalizability within URL research and capability in the socio-political domains:
  1. 1.
    Our framework integrates multi-view features into a single embedding, enabling the model to include in analysis user opinions and emotions, dynamic activity patterns, demographic profiles, and network interactions, making it better suited for tackling socio-political questions compared to existing models.
  2. 2.
    We cohesively integrate advanced methodologies into a unified framework, learning temporal behaviors even from sparse activity data using triplet representations, employing transformer architectures for embedding, and utilizing contrastive learning and network link prediction to capture user similarity more effectively.
  3. 3.
    Unlike task-specific approaches, our framework is universal and adaptable to a broad range of applications.
  4. 4.
    Our framework has demonstrated scalability, handling datasets with up to 17 million texts.
Our proposed framework helps to bridge the divide between user representation learning and socio-political analysis by leveraging features from dynamic and heterogenous users to improve classification. This work paves the way for more informed decisions and interventions. The source code is available at https://anonymous.4open.science/r/ts˙clustering-1C10.
Related Work
User representation learning has gained widespread interest within the recommender systems due to its ability to capture meaningful and compact embeddings of users’ behaviors and preferences (Yuan et al. 2021). Over the years, researchers have developed many methods, mostly for commercial applications (Li and Zhao 2021; Purificato et al. 2024). These include matrix and tensor factorization (Bhargava et al. 2015), auto-encoders (Zhuang et al. 2017), transformer-based architectures (Cheng et al. 2021; Hou et al. 2022; Shin et al. 2023), contrastive learning (Oord et al. 2018; Cheng et al. 2021), graph neural networks (Liu et al. 2023), and large language models (LLM) (Ren et al. 2024). Improving from task-specific methods (Guo et al. 2017), researchers have also explored universal user representation learning methods which can be generalized to different downstream tasks in recommendations (Yuan et al. 2021; Hou et al. 2022; Shin et al. 2023; Fazelnia et al. 2024).
Beyond recommendations, URL also presents considerable opportunities in the social and political domains, enabling a deeper understanding of public sentiment and societal trends (Pan and Ding 2019). However, the approaches designed for recommendations are built on commercially relevant features and/or pre-trained using commercial data (Li and Zhao 2021; Purificato et al. 2024), and thus are not compatible in social analyses. Moreover, most current approaches in social domains rely on a narrow subset of user features. For example, Mueen et al. (2016) and Nwala et al. (2023) uses temporal activity features to detect online bot and IO drivers; Hallac et al. (2021) experimented with different textual embedding methods; Perozzi et al. (2014) and Wang et al. (2017) uses network features and node prediction to learn user interests or communities. While multi-view studies are available, they are typically limited to integrating only text and network features, with many being tailored to specific tasks or platforms (Ribeiro et al. 2018; wan 2019; Lai and Neville 2020; Feng et al. 2021; Shen et al. 2023). There clearly lacks a universal framework for social analysis that is generalizable to different platforms and downstream tasks, as well as one that incorporates all of the textual, temporal, profile and network features into a multi-view embedding. Our framework, SoMeR, addresses these gaps.
Methods
@@
We propose a self-supervised framework to learn a latent user embedding space based on the architecture shown in Figure 1. From each user’s history, a timeline of texts, we first extract certain textual features, such as the sentence embeddings. Next, to better learn from users with sparse activities, we format the textual features and timestamps into triplets of observations $(timestamp,feature,value)$. These triplets pass through a Triplet Encoder, a transformer-based contextual learning module, and a fusion attention layer, being encoded into a user history embedding. We concatenate it with user’s profile embedding from a separate module, obtaining a complete user embedding. We train these encoding modules with two self-supervised objectives: network link prediction that learns patterns of interactions, including sharing, following or other connections, and contrastive loss that learns user similarity with respect to posting history. This method learns user similarity in a heterogeneous user population without the need for time-consuming human annotations. The method can be easily adapted for various downstream tasks such as supervised classification and unsupervised similarity search.
@@
In this study, we consider each user’s history to be a collection of timestamped texts. Each text can be an original post, a repost, a reply, etc. There can be other types of user activity, such as likes or views, that also provide valuable information that can be incorporated in future works. To learn from a user’s history, we first extract textual features that are useful for a given downstream task. In our experiments, we find that using contextual sentence-BERT embeddings (Reimers and Gurevych 2019) leads to powerful representation. However, BERT embeddings yield complex high-dimensional features, which can slow performance. To reduce model size and complexity, we perform dimensionality reduction using principal component analysis (PCA) to reduce BERT embeddings to the first five components and treat them as textual features. Prior works (Grootendorst 2022) have shown that meaningful contextual representation is still retained after dimension reduction operations on BERT embeddings.
@@
Social media users exhibit highly diverse posting behaviors. A small fraction of users generate most of the contents, whereas majority of users only have few posts over a long period of time (Krishnan et al. 2018). This leads to a sparse matrix of multivariate time series as the input data, making it hard to accurately learn an embedding space from sparse signals, and posing a challenge in time complexity. We take inspiration from (Tipirneni and Reddy 2022) and represent user posting history as a collection of triplets $(timestamp,feature,value)$. This only takes the observations available in the data, allowing us to account for the times with little to no activity. Thus, a dataset of $M$ users can be represented as $U$:
@@
Each user $u$ is characterized by their profile feature vector $d^{u}\in\mathbb{R}^{D}$ and their posting history $\mathbf{X}^{u}$, which is a set of $N$ triplets including the timestamp $t_{n}$, the feature category $f_{n}$ and the value of this feature category $v_{n}$. There can be more than one feature-value pairs at one time point, and therefore $N$ may be greater than the total number of time points.
Tipirneni and Reddy (2022) shows the effectiveness of using a feed-forward network to embed continuous values. We therefore use two separate feed-forward networks to encode timestamp $t_{n}$ and value $v_{n}$ into time embedding $e^{t}_{n}\in\mathbb{R}^{K}$ and value embedding $e^{v}_{n}\in\mathbb{R}^{K}$ respectively, where $K$ is the hidden dimension of these embeddings. These networks have one linear layer followed by a tanh activation function. For the feature categories, we use a lookup-table encoder to generate a feature embedding $e^{f}_{n}\in\mathbb{R}^{K}$. Lastly, we add up these three embeddings to be the triplet embedding $e_{n}^{triplet}\in\mathbb{R}^{K}$. Adding them up, rather than concatenating, has been proven its effectiveness (Tipirneni and Reddy 2022), and allows us to reduce dimensions of model weights to save training time and memory.
@@
The transformer architecture has been shown to have great performance in representation learning for time series and user behavior sequences (Zerveas et al. 2021; Shin et al. 2023). Therefore we choose to use the transformer encoder to better extract a latent representation from the triplet embedding $e_{n}^{triplet}\in\mathbb{R}^{K}$. We use $L$ transformer layers. Each layer has $H$ attention heads with learnable key, query and value weights $W^{k},W^{q},W^{v}\in\mathbb{R}^{K\times P}$ where $P$ is the hidden dimension size for these weights. After all attention heads are added up and layer normalized, the embedding vector is projected back to $K$-dimension through a feed-forward network. This network includes two layers with hidden dimension $2K$ and a ReLU activation in the middle. Lastly, layer normalization is applied again. From this transformer module we obtain $e_{n}^{trans}\in\mathbb{R}^{K}$. In our experiments and applications, data sizes are within 20 million, and therefore we choose to have a small transformer module with $L=2$, $H=4$ and $P=K/H$. We choose the hidden dimension $K=64$ with a grid search in $[32,64,128]$. We show in three applications that these parameters work well for our purposes.
@@
Individuals connected in social networks tend to be similar, and their behaviors are often affected by other users in the network (McPherson et al. 2001). Therefore it’s crucial to include network connection features when learning user representations (AlMahmoud and AlKhalifa 2018). We design a self-supervised network link prediction objective to train our model to learn interaction activities such as sharing, following and commenting. It is a feed-forward module to perform link prediction, with a binary cross-entropy loss. During training, we consider all pairs of distinct users in each batch. The feature for link prediction is the concatenated embedding of a user pair, and we obtain the links from a self-defined network as the binary labels (e.g. whether a user repost another user). This is described as following:
@@
Contrastive Learning
Contrastive learning aims to obtain a latent embedding space in which similar samples are closer and distinct samples are farther from each other. Many prior works in user representation learning have shown its success (Cheng et al. 2021; Shin et al. 2023). We adopt self-supervised contrastive learning to learn user similarity in their posting histories. The InfoNCE (Oord et al. 2018) loss function uses categorical cross-entropy loss to optimize the negative log probability of classifying one positive or similar sample correctly among a set of negative or unrelated samples. In our case, it can be written as:
@@
To generate the positive sample paired to each user history embedding, we perform data augmentation on users’ triplet data. Although researchers have proposed many time series data augmentation methods (Wen et al. 2020), many classical methods are not applicable to our scenario. For example, a user history cannot be sub-sequenced or shuffled in time domain. We use another efficient and simple way — bootstrapping with replacement from existing triplets $\{(t_{n},f_{n},v_{n})\}_{n=1}^{N}$. This ensures the generated set of triplets is similar. We incorporate noise into augmentation by
@@
During training, we tune $\gamma$ by grid searching in $[0.1,0.5,1,2,5]$. The results of the three applications below shows the effectiveness of this augmentation method.
@@
where $\lambda$ is a hyperparameter to balance between two losses. We perform grid search on the IO driver dataset for $\lambda$ in $[0.1,1,5,10]$ and find $\lambda=1$ gives the lowest validation loss. We use Adam optimizer, a learning rate of 5e-5, a cosine decay learning scheduler, an early stopping mechanism monitored by overall loss function, a maximum epoch of 60, and a batch size of 128. See Appendix Table 4 for a summary of all hyperparameters and their selecting procedures. We find changes to these hyperparamters do not strongly affect results.
@@
To verify that our self-supervised framework indeed learns both temporal activity and textual features from the triplet data, we test it on four synthetic datasets. We use synthesized numerical features to mimic textual features in real data. The four datasets simulate (1) a simple scenario with three clusters that are easier to distinguish, (2) a hard scenario with 10 clusters and more noise, (3) a scenario where clusters vary in temporal activity pattern but have the same features and values, and (4) a scenario where clusters vary only in features and values, but having all the same temporal patterns. On all datasets, our model can successfully detect the clusters, indicating it is able to learn heterogeneity in both temporal activities and numerical feature values (see Appendix).
In the next sections, we use three applications - IO driver detection, political polarization analysis, and hate group joiner prediction to illustrate the effectiveness and versatility of our framework.
Detecting Information Operation (IO) Drivers
Information operations use strategically organized efforts to manipulate public opinions at scale, contaminating the online information ecosystem with disinformation. The rapidly expanding social media platforms provide fertile grounds for these operations (Pacheco et al. 2021), necessitating their identification. IO drivers use different tactics from each other and from authentic accounts thus leading to distinct behaviors flagged by previous researchers, such as (1) co-sharing the same posts or URLs, (2) using an identical sequence of hashtags, (3) synchronized behaviors, and (4) sharing very similar posts (Pacheco et al. 2021; Nwala et al. 2023; Luceri et al. 2024). Our framework captures temporal, textual, profile, and network features, identifying platform-agnostic synchronized behaviors. We apply supervised fine-tuning on this framework and test it across three campaigns on X to demonstrate its effectiveness.
@@
We evaluate our method on an X dataset that prior IO driver detection methods have benchmarked on (Nwala et al. 2023; Luceri et al. 2024). These drivers of multiple information operations in 21 countries were suspended and released by X because they were associated with malicious IOs and violated the platform terms. We select four campaigns — one based in China involving a large number of accounts, one small operation in Egypt and UAE, and two operations in Venezuela which are combined into one dataset for testing the multi-campaign scenario.
We collect a set of control users by first collecting the top five keywords and top five hashtags for each IO driver within our X datasets. We then extract 10 random posts that were posted within the timeframe that the IO driver was active (between their first and last post). For each post, we find the post author and query all their posts made within the timeframe the IO driver is active. Table 1 shows the information about these campaigns. We randomly split data into 70% training, 15% validation and 15% testing.